OpenStudy (anonymous):

Find the exact value of the composite function: sinx ( tanx^-1 (-7/24)) Explain please!

OpenStudy (anonymous):



OpenStudy (anonymous):

there is a picture of an angle whose tangent is \(\frac{7}{24}\) what you need is the hypotenuse, which you find either by pythagoras \[h=\sqrt{7^2+24^2}\] or else by remembering the \(7:24:25\) right triangle

OpenStudy (anonymous):



OpenStudy (anonymous):

now you can see that the sine is \(\frac{7}{25}\) except we need to remember that the tangent was negative, and therefore you were in quadrant 4, which means the sine is negative as well

OpenStudy (anonymous):

if you have a question let me know, all steps i think are there

OpenStudy (anonymous):

Oy well, how do we get to the answer (5√41)/41 o;

OpenStudy (anonymous):

hmmm i think you do not in fact i know you do not

OpenStudy (anonymous):

maybe that was for another question

OpenStudy (noelgreco):

Tan is <0 in quadrants II and IV. If the question means explicitly \[\frac{ -7 }{ 24 }\] and not \[-\frac{ 7 }{ 24 }\] then we are truly confined to the 4th quadrant. Otherwise, the solution in II is also correct.

OpenStudy (anonymous):

Well, the entire question was (because maybe I set it up wrong): Given that f(x) = sin x, g(x) = cos x, and h(x) = tan x, find the exact value of the composite function. f(h^-1 (-(7/24)))

OpenStudy (anonymous):

maybe your question was \[\sin(\tan^{-1}(\frac{5}{6}))\]

OpenStudy (anonymous):

OMG.

OpenStudy (anonymous):

no you wrote it right

OpenStudy (anonymous):

NO WAIT. I SWITCHED THE QUESTIONS AROUND.

OpenStudy (anonymous):

was it \[\sin(\tan^{-1}(-\frac{5}{6}))\]

OpenStudy (anonymous):

the answer to that one would be \(-\frac{5}{\sqrt{41}}\)

OpenStudy (anonymous):

Wait, no I didn't

OpenStudy (anonymous):

That's how it was written

OpenStudy (anonymous):

@NoelGreco there is no difference between \(-\frac{7}{24}\) and \(\frac{-7}{24}\)

OpenStudy (anonymous):

and the answer was right that you had, I was just looking at the wrong one

OpenStudy (anonymous):

Sorryy o;

OpenStudy (anonymous):

whew !!

OpenStudy (anonymous):

is the method clear?

OpenStudy (anonymous):

Yes! (: Thank you very much!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

also please note that the domain of \(\tan^{-1}(x)\) is \([-\frac{\pi}{2},\frac{\pi}{2}]\) so you are in fact in quadrant 4, and not 2

OpenStudy (anonymous):

check that i mean the range, not the domain. the domain is all real numbers

OpenStudy (anonymous):

Mhm! (:

OpenStudy (anonymous):

the range of arctangent is \([-\frac{\pi}{2},\frac{\pi}{2}]\) so if your answer is negative you are in quadrant 4

OpenStudy (anonymous):

Alrighty o: Ty!

OpenStudy (anonymous):

yw (again)